# Speedy-Splat: Fast 3D Gaussian Splatting with Sparse Pixels and Sparse Primitives

### v2

@@ Abstract
Figure 1: We reduce the number of Gaussians by over $90\%$, only marginally decrease PSNR, and accelerate rendering speed by 6.2×6.2$\times$6.2 × in the Tanks & Temples truck scene when compared to 3D Gaussian Splatting (3D-GS). Additionally, we speed up training time by 1.38×1.38$\times$1.38 ×.
@@ 1 Introduction
3D-GS implements tile rendering to localize Gaussians in the image plane by assigning each Gaussian to the tiles that it intersects. We find the existing algorithm to be overly conservative; in Section 4.1.1, we address this by introducing our SnugBox algorithm to precisely localize Gaussians by computing a tight bounding box around their extent. Then, in Section  4.1.2, we extend SnugBox with our AccuTile algorithm to identify exact tile intersections. Both approaches are plug-and-play, lead to respective inference speed-ups of 1.82×1.82$\times$1.82 × and 1.99×1.$9$ on average, and do not change the 3D-GS renderings.
To reduce the total number of Gaussians while preserving visual fidelity, we extend an existing pruning method [11] by reducing its memory requirement by 36×36$\times$36 × and incorporating it into the 3D-GS training pipeline. During the densification stage of 3D-GS training, Gaussians are regularly replicated and pruned. In Section 4.2.2, we augment the densification stage with our Soft Pruning method to prune $80\%$ of Gaussians at regular intervals. After the densification stage, our Hard Pruning method, described in Section 4.2.3, prunes an additional $30\%$ of Gaussians at set intervals.
Our Speedy-Splat approach integrates both techniques into the 3D-GS training pipeline. On average, rendering speed is accelerated by 6.71×6.71$\times$6.71 ×, model size is reduced by 10.6×10.6$\times$10.6 ×, and training speed is improved by 1.4×1.4$\times$1.4 × across all evaluated scenes while maintaining high image quality.
@@ 2.1 Pruning
PUP 3D-GS [11] introduces a more principled approach by deriving a Hessian for each Gaussian that represents its sensitivity to the reconstruction error. While PUP 3D-GS achieves state-of-the-art post-hoc pruning results, computing its sensitivity score incurs considerable storage requirements that limit its viability for use during training. Our pruning approach directly extends PUP 3D-GS by improving its memory efficiency by 36×36$\times$36 ×.
@@ 3 Background
+SnugBox 0.656 0.046 0.208 (2.738$\times$) 0.729 (2.126$\times$) 0.041 (1.980$\times$) 2.344 (1.913$\times$) 4.102 (1.823$\times$)
+AccuTile 0.668 0.046 0.221 (2.575$\times$) 0.612 (2.533$\times$) 0.035 (2.326$\times$) 2.062 (2.175$\times$) 3.748 (1.995$\times$)
+Soft Pruning 0.370 (1.798$\times$) 0.030 (1.494$\times$) 0.146 (3.906$\times$) 0.404 (3.843$\times$) 0.024 (3.422$\times$) 1.337 (3.354$\times$) 2.381 (3.141$\times$)
+Hard Pruning 0.091 (7.293$\times$) 0.016 (2.769$\times$) 0.090 (6.325$\times$) 0.215 (7.217$\times$) 0.013 (6.537$\times$) 0.619 (7.247$\times$) 1.114 (6.712$\times$)
@@ 4.1.1 SnugBox
After the bounding box edges are identified, our method follows 3D-GS and converts these edges to tile indices by dividing by tile size, rounding, and clipping to the image boundary. As depicted in Figure 2, Snugbox can produce a significantly tighter bounding box than 3D-GS. Meanwhile, its computational overhead is small as it performs a constant number of operations and is only called twice in the rendering pipeline – once to count the Gaussians intersecting tiles in preprocess and once to populate the Gaussian-to-tile arrays in duplicateWithkeys. Table 1 reports that SnugBox improves the efficiency of all downstream functions and produces an average overall speed-up of 1.82×1.82$\times$1.82 ×.
@@ 4.1.2 AccuTile
   Initialize: Tile count C0absent0$\leftarrow$ 0← 0
  $\leftarrow$ Rb
  if  $\geq$ Bb then
@@
     $\leftarrow$ Bl if Bl in r else min(imin, imax)
     $\leftarrow$ Br if Br in r else max(imin, imax)
     tilemin, $\leftarrow$ Convert(emin, emax)
@@
Figure 2 illustrates how Accutile restricts the tight bounding box produced by SnugBox to the exact tiles that the Gaussian touches. The tile extent rows from Snugbox, shown in Figure 2(b), are processed starting from the bottom. The lower tile row boundary falls below the bounding box, so no initial intersection is computed. The upper boundary lies below the top of the bounding box, so intersection points $A$ and $B$ are calculated using Equation 15. $x_{\min}$ is assigned as this row’s minimum ellipse extent $e_{\min}$ because it lies within it, and $B$ is assigned as its maximum ellipse extent $e_{\max}$ because it is the maximal point. Consequently, this row’s tile extent is from the tile containing $x_{\min}$ to the tile containing $B$. For the next row, we keep points $A$ and $B$ and compute upper boundary points $C$ and $D$. $A$ and $x_{\max}$ are assigned as the row’s $e_{\min}$ and $e_{\max}$, and the tiles containing them are its tile extent. Finally, for the last row, $C$ and $D$ are kept; no additional points are computed because the row’s upper boundary is above the bounding box. Thus, the tile extent of this row is between $C$ and $D$.
The number of tiles is further reduced from Figure 2(b) to Figure 2(c), much less than the original 3D-GS method in Figure 2(a). In Table 1, we report that AccuTile further accelerates all downstream functions, culminating in an average overall speed-up of 1.99×1.$9$.
@@ 4.2.1 Efficient Pruning Score
Gradient $\nabla_{g_{i}}I_{\mathcal{G}}$ is already computed in the backward pass of render and can be efficiently squared and aggregated across all pixels. Moreover, the maximum space requirement for this score is proportional to the number of Gaussians $N$, reducing the storage requirement by 36×36$\times$36 × and allowing this score to be used during training.
@@ 4.2.3 Hard Pruning
We also observe that the model’s performance after densification closely matches that of the fully-trained model. The iterations after the densification stage essentially fine-tune the model and can be used to further “refine” it after pruning, similar to PUP 3D-GS [11] and LightGaussian [6]. In practice, our Hard Pruning method prunes the model by a constant ratio every $3000$ iterations starting at iteration $15000$. We Hard Prune $30\%$ of Gaussians in each interval, which, when paired with Soft Pruning, empirically reduces the total number of Gaussians across scenes by 10.6×10.6$\times$10.6 ×.
@@ 5.2 Implementation Details
+SnugBox 2.97M 244 (1.82$\times$) 21.2 (1.09$\times$)
+AccuTile 2.97M 267 (1.99$\times$) 21.0 (1.10$\times$)
+Soft Pruning 1.64M (1.79$\times$) 420 (3.14$\times$) 17.5 (1.32$\times$)
+Hard Pruning 0.28M (10.6$\times$) 898 (6.71$\times$) 15.7 (1.47$\times$)
@@
3D-GS [14] 1.00$\times$ 1.00$\times$ 1.00$\times$ 27.55 0.814 0.222
Trimming [2] 4.00$\times$ - - 27.13 0.798 0.248
Compact [18] 2.28$\times$ 1.07$\times$ 0.73$\times$ 27.08 0.798 0.247
EAGLES [10] 3.68$\times$ 1.51$\times$ 1.37$\times$ 26.94 0.800 0.250
Reducing [26] 2.33$\times$ 1.60$\times$ 1.23$\times$ 27.10 0.809 0.226
Light [6] 2.94$\times$ 1.76$\times$ - 27.28 0.805 0.243
ELMGS [1] 5.00$\times$ 2.69$\times$ - 27.00 0.779 0.286
PUP [11] 8.65$\times$ 2.55$\times$ - 26.83 0.792 0.268
Mini-Splat [7] 6.84$\times$ 3.20$\times$ 1.26$\times$ 27.34 0.822 0.217
+SnugBox 0.99$\times$ 1.81$\times$ 1.08$\times$ 27.55 0.814 0.221
+AccuTile 0.99$\times$ 1.99$\times$ 1.10$\times$ 27.57 0.814 0.221
+Soft Pruning 1.79$\times$ 3.14$\times$ 1.30$\times$ 27.32 0.807 0.246
+Hard Pruning 10.6$\times$ 6.51$\times$ 1.45$\times$ 26.94 0.782 0.296
@@ 5.3.1 Additive method performance
The efficacy of Speedy-Splat is demonstrated by Table 1, where we record the average execution times of each function across all scenes when additively applying our methods. SnugBox and AccuTile each introduce minimal additional computation time to preprocess and InclusiveSum. However, limiting the number of tiles touched accelerates all downstream functions, culminating in an overall speed-up of 1.82×1.82$\times$1.82 × by SnugBox that is raised to 1.99×1.$9$ by AccuTile. Applying soft pruning reduces the runtime of all functions by reducing the number of Gaussians, leading to a 3.14×3.14$\times$3.14 × overall speed-up. Finally, performing hard pruning improves overall speed by a whopping 6.71×6.71$\times$6.71 × over the baseline 3D-GS model.
@@ 5.3.2 Overall Performance
Table 3 compares our methods against other methods that reduce the number of Gaussians and increase inference speed using the mean of each metric across all scenes in the Mip-NeRF 360 dataset. The underlined methods are “lossless”, meaning that they avoid degrading visual fidelity at all. SnugBox and AccuTile, our lossless methods, improve rendering and training speed while leaving image quality metrics essentially unchanged or slightly better. Our full pipeline, labeled as “+Hard Pruning" boasts the highest compression ratios, rendering speeds, and training speed-ups across all datasets. Furthermore, its image quality metrics are competitive with the other methods.
@@ 7 Conclusion
In this work, we present Speedy-Splat: a new 3D-GS technique that accurately localizes Gaussians during rendering and significantly improves inference speed, model size, and training time. Enhanced localization is achieved by our SnugBox and AccuTile methods, while model size reduction is accomplished by our Soft and Hard Pruning approaches. Together, our Speedy-Splat methods accelerate rendering speed by an average of 6.71×6.71$\times$6.71 ×, reduce model size by 10.6×10.6$\times$10.6 ×, and improve training time by 1.47×1.47$\times$1.47 × across all scenes from the Mip-NeRF 360, Tanks & Temples, and Deep Blending datasets.
@@ A.1 AccuTile Proof of Correctness Sketch
Figure 7: We sweep pruning percentages in $5\%$ increments for Hard Pruning ($0\%-40\%$) and Soft Pruning ($0\%$, $50\%-95\%$) on all scenes listed in Section 5.1. Experiments are performed 3×3$\times$3 × on each scene without our Gaussian localization methods; the reported metrics are averaged across all runs. ($0\%$$0\%$) is the baseline 3D-GS model, the first column ($0\%$, :) is Hard Pruning in isolation, and the first row (:, $0\%$) is Soft Pruning in isolation. The red dots at ($80\%$, $30\%$) denote the percentage settings used in our manuscript. We report the FPS increase and the Number of Gaussians and Train Time decrease factors to be consistent with the format in Table 3.
@@
3D-GS [14] 1.00$\times$ 1.00$\times$ 1.00$\times$ 23.70 0.849 0.178
Trimming [2] 4.00$\times$ - - 23.69 0.831 0.210
Compact [18] 2.19$\times$ 1.16$\times$ 0.76$\times$ 23.32 0.831 0.201
EAGLES [10] - 1.73$\times$ 1.19$\times$ 23.10 0.820 0.220
Reducing [26] 2.56$\times$ 1.91$\times$ 1.27$\times$ 23.57 0.840 0.188
Light [6] 2.94$\times$ 1.97$\times$ - 23.11 0.817 0.231
ELMGS [1] 5.00$\times$ 4.05$\times$ - 23.90 0.825 0.233
PUP [11] 10.0$\times$ 4.00$\times$ - 22.72 0.801 0.244
Mini-Splat [7] 9.20$\times$ - - 23.18 0.835 0.202
+SnugBox 0.99$\times$ 1.61$\times$ 1.11$\times$ 23.69 0.849 0.178
+AccuTile 0.99$\times$ 1.67$\times$ 1.12$\times$ 23.73 0.849 0.177
+Soft Pruning 1.69$\times$ 2.48$\times$ 1.36$\times$ 23.54 0.841 0.201
+Hard Pruning 10.1$\times$ 6.30$\times$ 1.58$\times$ 23.45 0.821 0.241
@@ A.2 Overall Pruning Percent Metrics
In Figure 7, we perform a parameter sweep over Hard Pruning percentages from $0\%-40\%$ at $5\%$ intervals and Soft Pruning percentages at $0\%$ and from $50-95\%$ at $5\%$ intervals. We conduct each experiment 3×3$\times$3 × on each scene listed in Section 5.1 to reduce variance, then average the metrics across all runs. All experiments are run without our Gaussian localization methods – SnugBox and AccuTile – to ablate the effect of each pruning method in isolation. Our ($80\%$, $30\%$) pruning percentages are empirically selected to produce a favorable balance between speed and quality.
@@
3D-GS [14] 1.00$\times$ 1.00$\times$ 1.00$\times$ 29.09 0.886 0.288
Trimming [2] 1.33$\times$ - - 29.43 0.897 0.267
Compact [18] 2.65$\times$ 1.37$\times$ 0.79$\times$ 29.79 0.901 0.258
EAGLES [10] - 1.30$\times$ 1.31$\times$ 29.92 0.900 0.250
Reducing [26] 2.86$\times$ 1.79$\times$ 1.27$\times$ 29.63 0.902 0.249
@@
ELMGS [1] 5.00$\times$ 4.15$\times$ - 29.24 0.894 0.273
PUP [11] 10.0$\times$ 4.51$\times$ - 28.85 0.881 0.301
Mini-Splat [7] 8.06$\times$ - - 29.98 0.908 0.253
+SnugBox 0.97$\times$ 2.11$\times$ 1.12$\times$ 29.18 0.886 0.287
+AccuTile 0.97$\times$ 2.32$\times$ 1.13$\times$ 29.12 0.885 0.288
+Soft Pruning 1.86$\times$ 3.56$\times$ 1.41$\times$ 29.29 0.889 0.296
+Hard Pruning 11.1$\times$ 7.46$\times$ 1.57$\times$ 29.32 0.887 0.311
@@ A.5 Per-Scene Metrics
Baseline 71 201 142 126 164 91 117 172 140 141 200 138 185
+SnugBox 154 358 276 301 267 146 197 335 301 228 320 247 282
(2.15$\times$) (1.78$\times$) (1.95$\times$) (2.39$\times$) (1.62$\times$) (1.60$\times$) (1.68$\times$) (1.95$\times$) (2.15$\times$) (1.61$\times$) (1.60$\times$) (1.79$\times$) (1.53$\times$)
+AccuTile 168 413 330 332 285 155 221 378 315 248 343 272 294
(2.35$\times$) (2.05$\times$) (2.33$\times$) (2.64$\times$) (1.73$\times$) (1.70$\times$) (1.89$\times$) (2.20$\times$) (2.25$\times$) (1.75$\times$) (1.71$\times$) (1.97$\times$) (1.59$\times$)
+Soft Pruning 241 601 505 497 419 255 425 612 549 379 518 423 477
(3.37$\times$) (2.99$\times$) (3.56$\times$) (3.95$\times$) (2.55$\times$) (2.80$\times$) (3.63$\times$) (3.56$\times$) (3.92$\times$) (2.68$\times$) (2.59$\times$) (3.06$\times$) (2.58$\times$)
+Hard Pruning 662 978 842 1122 825 640 809 1277 942 724 1392 957 1149
(9.25$\times$) (4.87$\times$) (5.94$\times$) (8.91$\times$) (5.02$\times$) (7.03$\times$) (6.90$\times$) (7.42$\times$) (6.73$\times$) (5.12$\times$) (6.95$\times$) (6.93$\times$) (6.21$\times$)
@@
(1.13$\times$) (1.07$\times$) (1.11$\times$) (1.06$\times$) (1.08$\times$) (1.08$\times$) (1.11$\times$) (1.05$\times$) (1.08$\times$) (1.13$\times$) (1.09$\times$) (1.14$\times$) (1.09$\times$)
@@
(1.15$\times$) (1.08$\times$) (1.13$\times$) (1.07$\times$) (1.10$\times$) (1.09$\times$) (1.12$\times$) (1.06$\times$) (1.08$\times$) (1.14$\times$) (1.09$\times$) (1.15$\times$) (1.10$\times$)
@@
(1.38$\times$) (1.20$\times$) (1.30$\times$) (1.23$\times$) (1.39$\times$) (1.37$\times$) (1.30$\times$) (1.25$\times$) (1.27$\times$) (1.33$\times$) (1.38$\times$) (1.43$\times$) (1.37$\times$)
@@
(1.62$\times$) (1.28$\times$) (1.36$\times$) (1.38$\times$) (1.59$\times$) (1.49$\times$) (1.40$\times$) (1.41$\times$) (1.43$\times$) (1.55$\times$) (1.61$\times$) (1.62$\times$) (1.52$\times$)

### v3

@@ Abstract
Figure 1: We reduce the number of Gaussians by over $90\%$, only marginally decrease PSNR, and accelerate rendering speed by $6.2\times$ in the Tanks & Temples truck scene when compared to 3D Gaussian Splatting (3D-GS). Additionally, we speed up training time by $1.38\times$.
@@ 1 Introduction
3D-GS implements tile rendering to localize Gaussians in the image plane by assigning each Gaussian to the tiles that it intersects. We find the existing algorithm to be overly conservative; in Section 4.1.1, we address this by introducing our SnugBox algorithm to precisely localize Gaussians by computing a tight bounding box around their extent. Then, in Section  4.1.2, we extend SnugBox with our AccuTile algorithm to identify exact tile intersections. Both approaches are plug-and-play, lead to respective inference speed-ups of $1.82\times$ and $1.99\times$ on average, and do not change the 3D-GS renderings.
To reduce the total number of Gaussians while preserving visual fidelity, we extend an existing pruning method [11] by reducing its memory requirement by $36\times$ and incorporating it into the 3D-GS training pipeline. During the densification stage of 3D-GS training, Gaussians are regularly replicated and pruned. In Section 4.2.2, we augment the densification stage with our Soft Pruning method to prune $80\%$ of Gaussians at regular intervals. After the densification stage, our Hard Pruning method, described in Section 4.2.3, prunes an additional $30\%$ of Gaussians at set intervals.
Our Speedy-Splat approach integrates both techniques into the 3D-GS training pipeline. On average, rendering speed is accelerated by $6.71\times$, model size is reduced by $10.6\times$, and training speed is improved by $1.4\times$ across all evaluated scenes while maintaining high image quality.
@@ 2.1 Pruning
PUP 3D-GS [11] introduces a more principled approach by deriving a Hessian for each Gaussian that represents its sensitivity to the reconstruction error. While PUP 3D-GS achieves state-of-the-art post-hoc pruning results, computing its sensitivity score incurs considerable storage requirements that limit its viability for use during training. Our pruning approach directly extends PUP 3D-GS by improving its memory efficiency by $36\times$.
@@ 3 Background
[…]
@@ 4.1.1 SnugBox
After the bounding box edges are identified, our method follows 3D-GS and converts these edges to tile indices by dividing by tile size, rounding, and clipping to the image boundary. As depicted in Figure 2, Snugbox can produce a significantly tighter bounding box than 3D-GS. Meanwhile, its computational overhead is small as it performs a constant number of operations and is only called twice in the rendering pipeline – once to count the Gaussians intersecting tiles in preprocess and once to populate the Gaussian-to-tile arrays in duplicateWithkeys. Table 1 reports that SnugBox improves the efficiency of all downstream functions and produces an average overall speed-up of $1.82\times$.
@@ 4.1.2 AccuTile
[…]
@@
Figure 2 illustrates how Accutile restricts the tight bounding box produced by SnugBox to the exact tiles that the Gaussian touches. The tile extent rows from Snugbox, shown in Figure 2(b), are processed starting from the bottom. The lower tile row boundary falls below the bounding box, so no initial intersection is computed. The upper boundary lies below the top of the bounding box, so intersection points $A$ and $B$ are calculated using Equation 15. $x_{\min}$ is assigned as this row’s minimum ellipse extent $e_{\min}$ because it lies within it, and $B$ is assigned as its maximum ellipse extent $e_{\max}$ because it is the maximal point. Consequently, this row’s tile extent is from the tile containing $x_{\min}$ to the tile containing $B$. For the next row, we keep points $A$ and $B$ and compute upper boundary points $C$ and $D$. AA and $x_{\max}$ are assigned as the row’s $e_{\min}$ and $e_{\max}$, and the tiles containing them are its tile extent. Finally, for the last row, $C$ and $D$ are kept; no additional points are computed because the row’s upper boundary is above the bounding box. Thus, the tile extent of this row is between $C$ and $D$.
The number of tiles is further reduced from Figure 2(b) to Figure 2(c), much less than the original 3D-GS method in Figure 2(a). In Table 1, we report that AccuTile further accelerates all downstream functions, culminating in an average overall speed-up of $1.99\times$.
@@ 4.2.1 Efficient Pruning Score
Gradient $\nabla_{g_{i}}I_{\mathcal{G}}$ is already computed in the backward pass of render and can be efficiently squared and aggregated across all pixels. Moreover, the maximum space requirement for this score is proportional to the number of Gaussians $N$, reducing the storage requirement by $36\times$ and allowing this score to be used during training.
@@ 4.2.3 Hard Pruning
We also observe that the model’s performance after densification closely matches that of the fully-trained model. The iterations after the densification stage essentially fine-tune the model and can be used to further “refine” it after pruning, similar to PUP 3D-GS [11] and LightGaussian [6]. In practice, our Hard Pruning method prunes the model by a constant ratio every $3000$ iterations starting at iteration $15000$. We Hard Prune $30\%$ of Gaussians in each interval, which, when paired with Soft Pruning, empirically reduces the total number of Gaussians across scenes by $10.6\times$.
@@ 5.2 Implementation Details
[…]
@@
[…]
@@ 5.3.1 Additive method performance
The efficacy of Speedy-Splat is demonstrated by Table 1, where we record the average execution times of each function across all scenes when additively applying our methods. SnugBox and AccuTile each introduce minimal additional computation time to preprocess and InclusiveSum. However, limiting the number of tiles touched accelerates all downstream functions, culminating in an overall speed-up of $1.82\times$ by SnugBox that is raised to $1.99\times$ by AccuTile. Applying soft pruning reduces the runtime of all functions by reducing the number of Gaussians, leading to a $3.14\times$ overall speed-up. Finally, performing hard pruning improves overall speed by a whopping $6.71\times$ over the baseline 3D-GS model.
@@ 5.3.2 Overall Performance
Table 3 compares our methods against other methods that reduce the number of Gaussians and increase inference speed using the mean of each metric across all scenes in the Mip-NeRF 360 dataset. The underlined methods are “lossless”, meaning that they avoid degrading visual fidelity at all. SnugBox and AccuTile, our lossless methods, improve rendering and training speed while leaving image quality metrics essentially unchanged or slightly better. Our full pipeline, labeled as “+Hard Pruning” boasts the highest compression ratios, rendering speeds, and training speed-ups across all datasets. Furthermore, its image quality metrics are competitive with the other methods.
@@ 7 Conclusion
In this work, we present Speedy-Splat: a new 3D-GS technique that accurately localizes Gaussians during rendering and significantly improves inference speed, model size, and training time. Enhanced localization is achieved by our SnugBox and AccuTile methods, while model size reduction is accomplished by our Soft and Hard Pruning approaches. Together, our Speedy-Splat methods accelerate rendering speed by an average of $6.71\times$, reduce model size by $10.6\times$, and improve training time by $1.47\times$ across all scenes from the Mip-NeRF 360, Tanks & Temples, and Deep Blending datasets.
@@ A.1 AccuTile Proof of Correctness Sketch
Figure 7: We sweep pruning percentages in $5\%$ increments for Hard Pruning ($0\%-40\%$) and Soft Pruning ($0\%$, $50\%-95\%$) on all scenes listed in Section 5.1. Experiments are performed $3\times$ on each scene without our Gaussian localization methods; the reported metrics are averaged across all runs. ($\%00\%$) is the baseline 3D-GS model, the first column ($0\%$, :) is Hard Pruning in isolation, and the first row (:, $0\%$) is Soft Pruning in isolation. The red dots at ($80\%$, $30\%$) denote the percentage settings used in our manuscript. We report the FPS increase and the Number of Gaussians and Train Time decrease factors to be consistent with the format in Table 3.
@@
[…]
@@ A.2 Overall Pruning Percent Metrics
In Figure 7, we perform a parameter sweep over Hard Pruning percentages from $0\%-40\%$ at $5\%$ intervals and Soft Pruning percentages at $0\%$ and from $50-95\%$ at $5\%$ intervals. We conduct each experiment $3\times$ on each scene listed in Section 5.1 to reduce variance, then average the metrics across all runs. All experiments are run without our Gaussian localization methods – SnugBox and AccuTile – to ablate the effect of each pruning method in isolation. Our ($80\%$, $30\%$) pruning percentages are empirically selected to produce a favorable balance between speed and quality.
@@
[…]
EAGLES $0×\times 1.31$$\times$ 29.92 0.900 0.250
[…]
@@
[…]
@@ A.5 Per-Scene Metrics
Baseline 71 201 142 164 91 117 140 141 138 200 185 126 172
+SnugBox 154 358 276 267 146 197 301 228 247 320 282 301 335
[…]
+AccuTile 168 413 330 285 155 221 315 248 272 343 294 332 378
[…]
+Soft Pruning 241 601 505 419 255 425 549 379 423 518 477 497 612
[…]
+Hard Pruning 662 978 842 825 640 809 942 724 957 1392 1149 1122 1277
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]